OpenStudy (anonymous):

To help his sister, Lydia, pay for college, Jason offered her a loan at a simple annual interest rate of 2%. At the end of 4 years, Lydia paid off the loan, and Jason received $80 in interest. How much money did Jason loan his sister?

OpenStudy (shaik0124):

simple interest=(principal*time*rate)/100 as sinterest is $80 ,time is 4 years and rate is 2% now calculate principal that will be ur answer

OpenStudy (shaik0124):

@MinniMCMiner

OpenStudy (shaik0124):

$80=(p*4*2)/100 80*100=p*8 8000/8=p therefore p=$1000

OpenStudy (shaik0124):

jason offered $1000 to his sister

OpenStudy (shaik0124):

understood?